OpenStudy (anonymous):

what is the slope of the line passing throught the point (3, -4) and perpendicular to the line y=3x+5?

jimthompson5910 (jim_thompson5910):

The slope of y = 3x+5 is 3 or 3/1 Flip the fraction and flip the sign to get -1/3

jimthompson5910 (jim_thompson5910):

This is the slope of any line that is perpendicular to y = 3x+5

OpenStudy (anonymous):

right, so it would be 3?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

-1/3 because 3 is the original slope, -1/3 is the perpendicular slope

OpenStudy (anonymous):

the slope is -1/3?

jimthompson5910 (jim_thompson5910):

the slope of the perpendicular line is -1/3, yes

OpenStudy (anonymous):

Plug into the formula for the equation of a line: (y+4)=(-1/3)(x-3) = y+4=(-1/3x+1) y=-1/3x-3